A computer system has a 64 KB main memory and a 4 KB (data area only) cache. There are 8 bytes=cache line. Determine (1) the number of comparators needed and (2) the size of the tag field, for the following mapping scheme: Direct.
Answer #1

Given:

Main Memory size = 64KB = 2^6 KB = 2^6*2^{10} B = 2^{16} B

Cache size = 4KB = 2^2 KB = 2^2*2^{10} B = 2^{12} B

Cache line size = Block Size = 8 B = 2^{3} B

In Direct mapping,the address is divided into 3 fields. They are Tag bits, Line number, Block off-set represented as:

TAG LINE NUMBER BLOCK-OFFSET

From the given data we get,

Total number of bits in address = 16 (since main memory size = 2^{16} B)

Number of bits in block off-set = 3 (since block size = 2^{3} B)

Number of lines = \tiny \frac{Cache Size}{Block Size} = \tiny \frac{2^{12}}{2^3} = 2^9

Therefore, number of bits in line number = 9.

Finally, number of tag bits can be calculated as:

(Number of bits in main memory) - (Number of bits in block off-set + Number of bits in line number)

Number of tag bits = 16 - (9 + 3)

Tag bits = 16 - 12 = 4

(1) The number of comparators needed in direct mapping scheme = one

(2) The size of Tag field = 2^4 B = 16 B

Note:

1 KB = 2^{10} B